OpenStudy (anonymous):

y''-2y'+5y=5x^2-4x+2; y(0)=0 y'(0)=2

OpenStudy (turingtest):

got the complimentary solution yet?

OpenStudy (anonymous):

no =(

OpenStudy (turingtest):

that's just the solution to the homogeneous equation\[y''-2y'+5y=0\]can you do that?

OpenStudy (anonymous):

I can, I just don't know what to do with y(0)=0 e.c. given

OpenStudy (turingtest):

we are going to use that at the end to find the constants c1 and c2 do you have to use a particular method to find the particular solution? variation of parameters, undetermined coefficients, etc.?

OpenStudy (anonymous):

yeah, I need the particular solution

OpenStudy (turingtest):

I repeat: do you have to use a particular method to find the particular solution? variation of parameters, undetermined coefficients, etc.?

OpenStudy (anonymous):

I don't...

OpenStudy (turingtest):

Well I guess undetermined coefficients is what I jump to do you have any familiarity with that technique?

OpenStudy (anonymous):

I do

OpenStudy (turingtest):

ok, so you wanna try it on paper while I do the same?

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

ok I got my answer for the particular solution, let me know what you get so we can compare

OpenStudy (anonymous):

-x^2+6/5x-26\25

OpenStudy (turingtest):

er... no not quite what was your guess for the particular solution?

OpenStudy (anonymous):

well, f(x)=5x^2-4x+2=Ax^2+Bx+C F'(x)=2Ax+B F''(x)=2A 2A-2Ax-B-5Ax^2-5Bx-5C=5x^2-4x+2 A=-1 B=6/5 C=-26/25

OpenStudy (turingtest):

you just messed up on the system is all...

OpenStudy (turingtest):

wait, is it +5y or -5y ???

OpenStudy (turingtest):

...in the original problem I mean...

OpenStudy (turingtest):

also you forgot to multiply y' by 2...

OpenStudy (anonymous):

Oh... now i got it

OpenStudy (anonymous):

now A=1, B=-1,6, C=-0, 64 am I right?

OpenStudy (turingtest):

are you sure your original question is y''-2y'+5y=5x^2-4x+2 ??? (all the +'s and -'s correct?)

OpenStudy (turingtest):

and how do you have multiple answers for B and C o-0 ???

OpenStudy (anonymous):

yes, they are correct

OpenStudy (turingtest):

in that case you are still wrong, but closer A=1 is correct but how do you have multiple answers for the other constants?

OpenStudy (anonymous):

B=0!!!

OpenStudy (turingtest):

yeah ;)

OpenStudy (anonymous):

I've just found the mistake

OpenStudy (turingtest):

hooray! so your particular is...?

OpenStudy (anonymous):

it is x^2

OpenStudy (turingtest):

yep :) now the full solution is the linear combinations of the and complimentary solutions plus the particular \[y=y_c+y_p\]and what is that in your case?

OpenStudy (anonymous):

y=e^x(C1cos2x+C2sin2x)+x^2

OpenStudy (turingtest):

good, this is our \(general\) solution because it still has C1 and C2 in it to find the exact solution we need to apply the initial conditions we already know what y(x) is, so we can get an expression for y(0). what about y'(x) ? what then is y'(0) ?

OpenStudy (anonymous):

y(0)=C1=> C1=0

OpenStudy (turingtest):

yep :)

OpenStudy (turingtest):

and C2=?

OpenStudy (anonymous):

y'(x)= e^x(C2sin2x)+e^x*C2cosx*2+2x

OpenStudy (turingtest):

I think you had a slight typo on the cosine argument y'(x)= e^x(C2sin2x)+e^x*C2cos2x*2+2x

OpenStudy (turingtest):

but that doesn't change anything anyway :P

OpenStudy (turingtest):

so what is C2 ???

OpenStudy (anonymous):

C2 is 1

OpenStudy (turingtest):

right, so the final answer to out IVP is...?

OpenStudy (turingtest):

our*

OpenStudy (anonymous):

y=e^x*sin2x+x^2

OpenStudy (turingtest):

yes indeed :) and in case you don't believe me, ask the wolf http://www.wolframalpha.com/input/?i=solve+y%27%27-2y%27%2B5y%3D5x%5E2-4x%2B2+%2Cy%280%29%3D0%2C+y%27%280%29%3D2 (note that the wolf does not have the best reputation when it comes to solving DE's)

OpenStudy (anonymous):

Thank u so much!!!! Now I've made it out

OpenStudy (turingtest):

you're welcome :) see ya!

OpenStudy (anonymous):

I have 3 more equations to solve if you don't mind helping me